# On the application of Lorentz-Finsler geometry to model wave propagation

### v1

@@ 3.3.4 Step 4: The cut points
Nevertheless, the previous algorithm can be difficult to implement and expensive in computing time and power. An alternative option is to take advantage of Theorem 2 and choose $\tau$ small enough so that there are no intersections between wave trajectories in the discrete wavemap. Even though this time can be very small in theory, we can always choose carefully the initial points to avoid early intersections (e.g. choosing fewer initial points in regions where the wave trajectories tend to converge), at the cost of losing precision. In this case, =$\mathcal{E}=\{1,\ldots,m\}$.
@@ 3.3.5 Step 5: The final wavefront
Once we have computed $\mathcal{W}_{\tau}$, we can return to step 1 using $\mathcal{W}_{\tau}$ as the initial wavefront, and $t=\tau$ as the initial time, in order to find the wavefront at later times.
@@ 3.4 Advantages
  • Efficiency: the spirit of the model is to obtain the evolution of the wave by computing individual wave trajectories. This simplifies the direct approach of the classical wave equation and its generalizations—e.g. Richards’ equations for an elliptical propagation (see [24])—, where the wavefront is directly characterized and computed by means of a partial differential equation (PDE) system. Therefore, this model essentially reduces the general PDE system provided by Huygens’ principle (see [14, Theorem 4.14]) to the computation of several ODE ones—the geodesic equations—, which are easier to solve and more efficient, computationally speaking. Moreover, this approach also simplifies the management of cut points and possible crossovers of the wave: the PDE system cannot be solved beyond cut points (see [14, Remark 4.15]) and the wavefront has to be corrected each time one appears, whereas the individual wave trajectories can all be computed until the selected final time, regardless of the number of intermediate cut points, and then a single algorithm such as the one described in § 3.3.4 can be implemented to correct the wavefront (see also the discussion in [15, Appendix A.1]).
  • Accuracy: the model can handle any propagation pattern and any time, position and direction dependence, so long the infinitesimal spread (i.e. the indicatrix $\Sigma_{t,x}$) remains strongly convex. This is a major improvement over models that need to impose some restrictions on the wave spread, either because the differential equations they solve are only valid for a specific propagation shape—e.g. current fire growth simulators based on Richards’ equations need to assume that the infinitesimal fire spread always takes the form of an ellipse (see, e.g. [9, 27])—, or because this simplification is crucial for the model to be computationally efficient. In our model, every modification of the wave propagation is encoded within the (time-dependent) Finsler metric $F$ and the geodesic equations (along with the method used to solve them) are formally the same, regardless of how intricate $F$ is—i.e. increasing the complexity of $F$ does not entail a technical complication nor does it make the model inefficient.

### v2

@@ 3.3.4 Step 4: The cut points
Nevertheless, the previous algorithm can be difficult to implement and expensive in computing time and power. An alternative option is to take advantage of Theorem 2 and choose $\tau$ small enough so that there are no intersections between wave trajectories in the discrete wavemap. Even though this time can be very small in theory, we can always choose carefully the initial points to avoid early intersections (e.g., choosing fewer initial points in regions where the wave trajectories tend to converge), at the cost of losing precision. In this case, $\mathcal{E}=\{1,\ldots,m\}$.
@@ 3.3.5 Step 5: The final wavefront
Once we have computed $\mathcal{W}_{\tau}$, we can return to step 1 using $\mathcal{W}_{\tau}$ as the initial wavefront and $t=\tau$ as the initial time, in order to find the wavefront at later times.
@@ 3.4 Advantages
  • Efficiency: the spirit of the model is to obtain the evolution of the wave by computing individual wave trajectories. This simplifies the direct approach of the classical wave equation and its generalizations—e.g., Richards’ equations for an elliptical propagation (see [24])—where the wavefront is directly characterized and computed by means of a partial differential equation (PDE) system. Therefore, this model essentially reduces the general PDE system provided by Huygens’ principle (see [14, Theorem 4.14]) to the computation of several ODE ones—the geodesic equations—which are easier to solve and more efficient, computationally speaking. Moreover, this approach also simplifies the management of cut points and possible crossovers of the wave: the PDE system cannot be solved beyond cut points (see [14, Remark 4.15]) and the wavefront has to be corrected each time one appears, whereas the individual wave trajectories can all be computed until the selected final time, regardless of the number of intermediate cut points, and then a single algorithm such as the one described in § 3.3.4 can be implemented to correct the wavefront (see also the discussion in [15, Appendix A.1]).
  • Accuracy: the model can handle any propagation pattern and any time, position and direction dependence, so long as the infinitesimal spread (i.e., the indicatrix $\Sigma_{t,x}$) remains strongly convex. This is a major improvement over models that need to impose some restrictions on the wave spread, either because the differential equations they solve are only valid for a specific propagation shape—e.g., current fire growth simulators based on Richards’ equations need to assume that the infinitesimal fire spread always takes the form of an ellipse (see e.g. [9, 27])—or because this simplification is crucial for the model to be computationally efficient. In our model, every modification of the wave propagation is encoded within the (time-dependent) Finsler metric $F$ and the geodesic equations (along with the method used to solve them) are formally the same, regardless of how intricate $F$ is—i.e., increasing the complexity of $F$ does not entail a technical complication nor does it make the model inefficient.